Fermi National Laboratory


Accelerator Update for June 13, 2005 - June 15, 2005

Monday June 13
The day shift began with Operations monitoring the Antiproton Source (Pbar) stack, with NuMI off due to a ground fault and the Tevatron (TeV) off due to a CHL power failure, and with access to the Main Injector (MI), TeV, and Switchyard (SY) for maintenance work.

Linac came out of access and operators began turning the machine back on at 11:37 AM.

Operations reported at 11:56 AM that NuMI experts had found a bad cable on E:V100 in the MI line. There was no ground fault.

Operations established beam to the Booster dump at 1:46 PM.

Pbar experts began a study period at 5:29 PM.

Around 6:00 PM, Operations began troubleshooting a problem with the Linac Buncher. Beam resumed around 6:30 PM.

Operations established beam to MiniBooNE and NuMI at 6:32 PM.

A Recycler E-cool expert began a study at 6:40 PM.

At 6:59 PM, the NuMI beam permit dropped due to the current for the V100 power supply drifting out of tolerance. An operator called in an expert.

A power supply expert repaired V100. NuMI beam resumed at 8:22 PM.

A Pbar expert began a reverse proton study at 8:22 PM.

At 9:21 PM, the NuMI beam permit tripped off due to an E:V100 high temperature alarm. Operations reset it and called in an expert.

Operations began preparing MI for access 10:32 PM.

A NuMI expert reported at 11:33 PM from the MI tunnel that E:V100 had its LCW valved off. He valved the magnet in.

Tuesday June 14
The midnight shift began with Operations monitoring the stack, with operators making MI ready for beam, and with the TeV cryo system recovering from the CHL power outage.

Operations established beam to MiniBooNE at 12:45 AM.

At 12:48 AM, Operations reported that E:V100 was still tripping NuMI off. An operator called an expert and then, per instructions, changed the trip window limits. NuMI beam resumed at 12:56 AM.

At 1:23 AM, NuMI tripped off and then tripped again at 1:42 AM. It reset okay both times.

At 1:48 AM, a Booster kicker (MKS02) tripped off. Operators investigated and noticed it was sparking. They bypassed the kicker and called in an expert.

E-cool experts finished their studies at 1:55 AM.

At 2:02 AM, NuMI tripped off. Operations reset E:V100.

The Pbar studies ended at 2:19 AM, but Pbar couldn’t resume stacking due to Booster problems.

At 2:42 AM, NuMI tripped again. Per expert instructions, operators changed the trip limits.

The Booster expert reported at 3:42 AM that the tube for MKS02 had failed. He began its change out.

At 4:05 AM, NuMI tripped off for the first of three times in ninety minutes. It reset okay each time.

At 4:09 AM, Operations reported that the Pbar stack was lost, possibly due to an Accumulator quadrupole (A:LQ) power supply. Operators investigated and then called in an expert.

The Pbar power supply expert reported at 4:55 AM that the 480 volt circuit breaker for A:LQ was bad. He began repairs.

At 5:17 AM, NuMI tripped off due to E:V118. It reset okay.

Operations began preparing the Pbar tunnel for access at 5:24 AM.

The Booster expert reported at 5:25 AM that he had finished his repair of MKS02.

Operations reported at 8:26 AM that feeders 38 and 47 had been reenergized.

The Pbar power supply expert reported at 9:57 AM that he had replaced the breaker for A:LQ.

At 10:26 AM, NuMI tripped off twice in forty-five minutes. With permission, it was reset both times.

Operators began preparing MI for access at 2:57 PM to repair a door. Techs repaired the door and MI was ramped at 4:29 PM.

Operations tried sending beam to MiniBooNE at 4:37 PM, but problems with HV703 and V108 delayed the turn on. Experts began repairs around 6 PM.

Pbar experts began setting up for a reverse proton study at 6:07 PM.

At 6:28 PM, Operations reported vacuum problems in Pbar. Reverse proton studies ended at 9:48 PM and operators prepared the tunnel for access. The vacuum techs fixed the problem and were out of the tunnel by 10:38 PM.

Pbar experts resumed their reverse proton study at 11:25 PM.

Operation resumed sending beam to NuMI at 11:35 PM.

Wednesday June 15
The midnight shift began with the TeV and Debuncher cooldown continuing, with Pbar conducting reverse proton studies, and with MiniBooNE and NuMI taking beam.

At 3:01 AM, NuMI tripped off for the first of seven times over five hours due to beam positions. It reset okay each time.

At 4:36 AM, Operations reported that three Booster experts were coming in to help operators tune Booster.

At 5:48 AM, an LCW leak was noticed coming from the supply for a Booster RF station (BRF7). The station was turned off and an operator called in an expert.

Operations reported at 7:08 AM that the Pbar Debuncher cryo system was ready for normal operation.

Pbar studies finished at 7:11 AM.

The Plan for Wednesday
The plan for today is to finish work in the TeV. The TeV cryo system should be ready to go by midnight. TeV experts will conduct about eight hours of orbit studies once the cryo system is ready. A new store will go into the TeV on Thursday morning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/15/2005   email Fermilab